# Enabling decision support over confidential data

### v1

@@ Abstract
Enabling automated decision-making processes by leveraging data-driven analysis is a core goal of Decision Support Systems (DSSs). In multi-party scenarios where decisions rely on distributed and sensitive data, though, ensuring confidentiality, verifiability, transparency, integrity, and consistency at once remains an open challenge for DSSs. To tackle this multi-faceted problem, we propose the Secure Platform for Automated decision Rules via Trusted Applications (SPARTA) approach. By leveraging Trusted Execution Environments (TEEs) at its core, SPARTA ensures that the decision logic and the data remain protected. To guarantee transparency and consistency of the decision process, SPARTA encodes decision rules into verifiable software objects deployed within TEEs. To maintain the confidentiality of the outcomes while keeping the information integrity, SPARTA employs cryptography techniques on notarized data based on user-definable access policies. Based on experiments conducted on public benchmarks and synthetic data, we find our approach to be practically applicable and scalable.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.[23||24]xxxx www.ndss-symposium.org
I Introduction
Decision Support System (DSS) is an information system designed to help decision-makers solve complex problems and optimize decision processes. DSS enhances decision-making by providing data-driven insights, analytical tools, and structured problem-solving support. These systems facilitate efficient decision-making by leveraging IT-based solutions to process large volumes of information, evaluate multiple scenarios, and improve the overall effectiveness of strategic and operational decisions [1, 2]. They play a pivotal role in several fields including healthcare [3, 4], finance [5], business [6], and supply-chain [7] due to their ability to process large volumes of data, generate insights, and enhance decision accuracy. By efficiently evaluating various conditions and scenarios, these systems contribute to improved time and resource management while ensuring robust performance assessments. Ultimately, DSS enhances decision-making by offering reliable, adaptable, and user-centric solutions, even when problem parameters are not well-defined [8]. In multi-party processes, the decision-making support spans across multiple sources of information. In these cases, it is not easy to establish a trusted exchange of data and a common understanding of both the information and the decision logic [9, 10]. To be applicable in sensitive domains such as healthcare or finance, decision support systems must ensure the confidentiality and traceability of data, support user-definable decision logic, and guarantee system security and availability. Classical decision support systems that cater for customization support languages like Decision Model and Notation (DMN) [11], but are typically designed as modules of centralized information systems that do not guarantee privacy or traceability. Custom privacy-oriented approaches have been proposed to safeguard confidentiality and security but cannot provide means for user-definable decision logic or availability [12, 13] Blockchain-based solutions allow for customization and security and availability, but not for confidentiality [14].
In this paper, we present Secure Platform for Automated decision Rules via Trusted Applications (SPARTA), a novel solution that leverages Trusted Execution Environment (TEE) [15] to securely execute automatic decision support systems in multi-party processes. Our TEE-based solution enables automated decision-making support securely and privately while maintaining trust, transparency, and auditability.
To summarize, our main contributions are as follows:
@@ I Introduction
    We provide an open-source implementation of our solution, which we use to measure its performance, evaluate its scalability, and assess the overhead introduced by our secure platform with public, cross-domain benchmarks.
In the remainder of the paper, Sect. II outlines the core concepts we used to devise our solution. Sect. III presents a running example that is a reference throughout the paper and outlines the problem we tackle along with the requirements it poses. In Sect. IV, we describe our approach in detail. Sect. V provides a formal security analysis of our solution. In Sect. VI, we present our proof-of-concept implementation and the results of the experiments we conducted therewith. In Sect. VII, we discuss the implications and limitations of our work. Sect. VIII reviews related work. Finally, Sect. IX concludes the paper and outlines directions for future work. Following the open science principles, the source code and experiment data can be found at: anonymous.4open.science/r/SPARTA.
II Background
@@
Confidential computing represents a paradigm wherein sensitive data undergoes processing within a protected processing unit. Within this unit, the data and the processes treating those data remain concealed from any external entities [16]. Confidential computing underpins the Trusted Execution Environments (TEEs) [15]. A TEE is an environment that provides a level of assurance of the following three properties: (i) data confidentiality, namely the fact that unauthorized entities cannot access or observe data while it resides within the TEE and is being processed, (ii) data integrity, namely the fact that unauthorized entities cannot modify, delete, or insert data while it remains within the TEE and is undergoing processing and (iii) code integrity, namely that unauthorized entities cannot manipulate, delete, or introduce alterations to the code executed within the TEE [17, 18, 19]. Our work leverages Intel Software Guard Extensions (SGX) [20, 21, 22, 23], a collection of security-related instructions embedded within contemporary Intel Central Processing Unit (CPU)s. SGX facilitates the secure execution of code (deployed in the form of so-called Trusted Applications, or Trusted Apps for short) and data within enclaves, namely hardware-protected environments. The three primary security features of enclaves are isolation, attestation, and sealing. Isolation involves the enclave’s placement within a hardware-secured memory enclave page cache, specifically designated for housing protected code and data. However, TEE-allocated memory is limited. Exceeding these limits may entail stops that can severely affect availability (the service is interrupted) or data secrecy (encrypted data can be swapped to outer memory regions). Additionally, SGX enables remote attestation to allow external entities to authenticate a target enclave. During remote attestation, the external entity communicates with the attestation service to validate the integrity of the information provided by the target enclave. This process entails comparing enclave information with the trusted configuration within the target enclave, ensuring the integrity of code and data loading during the initial enclave establishment. Moreover, SGX employs sealing mechanisms to safeguard enclave contents stored beyond the enclave’s boundaries. These mechanisms entail encrypting enclave contents using a concealed sealing key, derivable from either the measurement hash or the author identity of the enclave. Both sealing mechanisms perform Advanced Encryption Standard - Galois/Counter Mode (AES-GCM) encryption. This approach guarantees exclusive access and decryption privileges to the sole corresponding enclave.
Decision logic and access control
One of the most well-known languages to model and formalize decision logic in business applications is Decision Model and Notation (DMN[24, 25, 26], a standard by the Object Management Group (OMG) 111omg.org/spec/DMN, accessed 2025-06-17 for modeling and formalizing decision logic in multi-party processes and business applications. A DMN, at its core, is a decision table that captures critical knowledge and automates decision logic [27]. A DMN table consists of columns representing the inputs and outputs of a decision, and rows denoting rules (the decision logic). Rules are expressed using Friendly Enough Expression Language (FEEL), a language defined within the DMN standard. Each row has an identifier, a condition for each input column, and one specific value for each output column. A rule is considered applicable when the input values satisfy all associated conditions; in that case, the corresponding output values are produced [27, 26]. To express authorization policies over the usage of resources, the Abbreviated Language for Authorization (ALFA) 222alfa.guide/alfa-authorization-language/, accessed 2025-06-17 language provides a concise, human-readable format. Built on top of the eXtensible Access Control Markup Language (XACML),333docs.oasis-open.org/xacml, accessed 2025-06-17 ALFA offers a simplified syntax for defining fine-grained, attribute-based access control (ABAC) policies based on users, actions, resources, and context.
Distributed notarization and storage
A blockchain is a Distributed Ledger Technology (DLT) wherein transactions are sorted, organized into blocks, and interconnected to form a chain. Blockchains possess resistance against tampering, achieved through cryptographic methods. Public blockchains like Ethereum [28] and Algorand [29] permit the deployment and execution of programs named smart contracts [30, 31]. To mitigate the expenses associated with the invocation of smart contracts, external Peer-to-peer (P2P) systems are frequently employed for storing large volumes of data [32]. Among the facilitating technologies is the InterPlanetary File System (IPFS) [33],444 ipfs.tech, accessed 2025-07-30 a distributed system designed for file storage and access, leveraging a Distributed Hash Table (DHT) to distribute files across multiple nodes. Utilizing content-addressing, IPFS assigns a unique identifier to each file within the network. When a file is uploaded to IPFS, it generates a content-addressed resource locator based on its hash (e.g., QmbPg[…]shTg9). Changing the file changes the hash and hence makes the locator invalid. IPFS supports mutable pointers that allow content updates without changing the shared reference. Specifically, InterPlanetary Name System (IPNS) enables the creation of mutable pointers, known as IPNS names (e.g., PatientsWAggrKey: QmbPg[…]shTg9). These names point to the content-addressed locators, and can be updated with newer locators while retaining the verifiability of content addressing.
III Problem illustration, running example, and requirements
@@
Table I: An excerpt of the data fields provided to the system by the process participants of Sect. III
@@
To illustrate the problem we address, we introduce a real-world scenario from the healthcare domain: a vaccine distribution campaign inspired by [34, 35] and linked to a research project funded by the European NextGenerationEU. [^5] To comply with the double-blindness requirement of the conference, we will reveal the project only upon acceptance. We use this scenario as a running example and motivating case study throughout the paper.
In our scenario, a central medical hub (henceforth, medical hub for short) receives vaccine supplies and handles their dispatch in a federal State. The downstream process involves other participants, namely patients who get an injection, vaccination centers who administer the injections, and several specialized carriers delivering the vaccines on premise. It covers two stages.
In the first phase, the involved participants act as Data Providers by contributing data into the system. We will refer to it as the data provision stage. In our scenario, e.g., the medical hub reports received vaccine stocks; patients submit vaccine requests including personal information; vaccination centers update storage capacity and vaccination progress; and carriers register their service conditions. Table I contains part of the data these participants must provide to the system.
In a subsequent stage (data usage), authorized participants retrieve and read the data provided earlier, either specific to single entities or aggregated, and receive decision support depending on their role in the process. For example, patients can retrieve their declarations at any time, carriers can be informed about the on-time delivery rate, and the medical hub requests automated support for vaccine supply distribution. Decisions related to the vaccine distribution include (i) the priority to be assigned to patients to administer the vaccine dose (which we will use as a running example throughout this paper), (ii) the restocking need for each vaccination center, (iii) and which carrier should handle each delivery.
In our example, several critical aspects need to be considered. First, all data is confidential and must be protected from unauthorized access. For example, revealing a single patient’s metadata constitutes a privacy breach. Second, data integrity and availability are critical. Any unauthorized modification could alter decisions or falsify patient records, while system failures or data unavailability could disrupt the decision-making process. Ensuring confidentiality, integrity, and availability while enabling a secure and verifiable automated decision-making is an open challenge motivating our research.
@@
IV Approach
In this section, we present our approach named Secure Platform for Automated decision Rules via Trusted Applications (SPARTA). We first introduce the system components (Sect. IV-A), and then describe how they interact to manage information and enable secure decision-making over confidential data (Sect. IV-B).
IV-A Actors and basic components
@@
Figure 1 provides a bird’s eye view of the SPARTA approach. We identify four main actors, each responsible for a distinct functionality. The Policymaker defines the decision rules and specifies which functions are available, along with the participant attributes or attribute combinations required to invoke them. Considering the example in Sect. III, the Policymaker role can be fulfilled either by a regulatory body such as the Ministry of Health, or by the consortium directly involved in the decision process (upon consensus, as in [36]). The Data Provider provides the input data and contextual information used for decision-making. In our example, the Data Providers are the participants: the medical hub, patients, vaccination centers, and the carriers. The Certification Authority (CA) signs a Certificate Signing Request (CSR) generated by a Registration Authority (RA), thereby producing a valid certificate. The RA handles participant registration and verifies their identity and role. Once verification is complete, it generates the CSR. By signing the CSR, the CA attests actor metadata, authenticates participants, and certifies their roles and attributes, determining their eligibility to interact with the system and invoke specific functions. In our example, the role of the CA should be fulfilled by a globally recognized Certification Authority (e.g., DigiCert, Entrust, or GlobalSign), [^6] DigiCert: digicert.com; Entrust: entrust.com; GlobalSign: globalsign.com/en, accessed 2025-08-04 akin to what is offered by the National Health Service in the United Kingdom [37]. The RA role should be covered by a National Health Service Provider or a government eID system for patients; a Regional Department of Public Health or Healthcare Ministry for vaccination centers; and a Ministry of Transport or logistics regulator for carriers. Finally, the Decider receives decision support based on the available data. If permitted by the access policies, the Decider may also retrieve raw data and invoke aggregation functions. In our example, the Decider is the medical hub.
@@
The interplay of five basic components realizes the architecture. (I) The Confidential Computing Unit (CCU) is SPARTA’s core component. We deploy the CCU inside a Trusted Execution Environment (TEE). This hardware-secured environment guarantees data confidentiality and ensures the integrity of both computation and code (see Sect. II), thereby meeting Tab. II. Notice that we make CCU replicable for the sake of information availability and system robustness (Tabs. II and II). The CCU deploys in a dedicated enclave within the TEE a separate Decision Trusted App (DeTApp) for each decision process. A DeTApp consists of the following. During SPARTA’s execution phase, it exposes (i) an invocable Decision Support Object (DeSObj) providing and controlling access to the Policymaker’s decision-making functions. The DeSObj, in turn, calls (ii) a Decision Function Object (DeFObj) realizing those functions, and a linked (iii) Decision Library (DeLib), a predefined library that offers utilities for DeSObjs such as data retrieval and decryption, policy checking, and certificate validation. The first two software objects are customizable and are built in SPARTA’s initialization phase by a DeSObj builder and a DeFObj builder, respectively, based on a policy descriptor and a decision model provided by the Policymaker. A DeTApp builder binds those objects together with the DeLib and deploys the DeTApp in the enclave.
@@
(III) The Admin Interface handles information exchange with CA and Policymaker. (IV) The User Interface is the means through which the Data Providers and the Decider interact with the CCU. Finally, SPARTA is paired with (V) the Data Manager, an external component providing two architectural buttresses: (i) The Data File Processor, storing the content of the exchanged data, and (ii) The Notarization System, saving and sealing the resource locators to those data. To meet Tab. II, we leverage two technologies for the realization of the above components. We employ IPNS over IPFS for the Data File Processor because IPNS guarantees permanence, immutability and content-addressing for stored files (see Sect. II). We designate a public blockchain platform as the Notarization System to store the IPNS names in a tamper-proof, available and traceable way. We remark that other solutions could be used such as STORJ777storj.io, accessed 2025-07-30 and distributed relational databases [38], respectively, in scenarios with less restrictive security requirements.
Next, we elucidate the interplay of the above components.
IV-B Workflow
The workflow of SPARTA mainly consists of two phases, as illustrated in Fig. 1: First, the system initialization (Sect. IV-B1) prepares the necessary data structures and programs, then the execution follows (Sect. IV-B2).
IV-B1 System initialization

### v2

# A TEE-based Approach for Security and Privacy in Decision Support

@@ Abstract
Decision Support Systems are increasingly adopted to automate decision-making processes across industries, organizations and governments. However, decision support requires maintaining data privacy, integrity, and availability while ensuring customization, security, and verifiability of the decision process. Existing solutions fail to guarantee those properties altogether. Most commercial tools cater for data integrity and process customization but are centralized. This centralization potentially compromises data privacy and availability, as well as process security and verifiability. To overcome these limitations, we propose SPARTA, an approach based on Trusted Execution Environments (TEEs) that automates decision processes. To maintain data privacy, integrity, and availability, SPARTA employs efficient cryptographic techniques on notarized data with access mediated through user-defined access policies. Our solution also allows users to define decision rules, which are translated to certified software objects deployed within TEEs, thereby guaranteeing customization, verifiability, and security of the process. Based on experiments conducted on public benchmarks and synthetic data, we show that our approach is scalable and adds limited overhead compared to non-cryptographically secured solutions.
1 Introduction
A Decision Support System (DSS) is an information system designed to help decision-makers solve complex problems and optimize decision processes. DSS s enhance decision-making by providing data-driven insights, analytical tools, and structured problem-solving support [ARNOTT2008657, liu2010integration]. They play a pivotal role in several fields including healthcare [NIU2024114228, BARRACOSA2023114046], finance [KWON2024114100], business [ARTELT2024114249], and supply-chain [PEREIRA2022113795] due to their ability to process large volumes of data, generate insights, and enhance decision accuracy. In cooperative systems, the decision-making support spans across multiple parties and sources of information, thus making it harder to establish a trusted data exchange and a common understanding of the decision logic [bazhenova2019bpmn, van2001p2p].
No current solution simultaneously guarantees data privacy, integrity and availability, while ensuring customizability, security and verifiability of decision processes. Systems that cater for decision process customization are typically designed as modules of centralized information systems that do not guarantee privacy or verifiability. Other solutions rely on cloud platforms, exposing all data and decision logic to third parties with the risk of potential leakage, manipulation, and unverifiable outcomes [LIU2018825]. Custom privacy-oriented approaches have been proposed to safeguard confidentiality and integrity by leveraging encryption mechanisms. Those solutions, however, rely on trusted authorities that handle the encryption keys, thereby introducing single points of failure or substantial overhead. Moreover, they typically fail to provide means for user-definable decision logic. Blockchain-based solutions have been proposed to allow for decision logic customization, security and availability employing smart contracts as automated decision makers. However, to enable validation, they publicly expose input data and its processing. Thus, privacy is compromised [haarmann2018dmn].
Against this background, we present a Secure Platform for Automated decision Rules via Trusted Applications (SPARTA), an approach that leverages Trusted Execution Environments (TEEs) [sabt2015trusted] to securely execute automatic decision support systems in multi-party cooperation scenarios. Our solution enables automated decision support securely and privately while maintaining trust, transparency, and auditability. We employ the TEE to to securely process encrypted data stored in a distributed file system (InterPlanetary File System, IPFS),4 to make data-driven decisions based on custom decision rules, and to use a public blockchain as a notarization system to guarantee information integrity and enable authorized users to access the data. We resort to well-established standards for specifying custom decision logic (Decision Model and Notation, DMN [Dumas.etal/2018:FundamentalsofBPM]) and for defining access policies (Abbreviated Language for Authorization, ALFA).2
Specifically, we provide the following contributions:
@@ 1 Introduction
    Using our open-source implementation of our solution, we measure its performance, evaluate its scalability, and assess the overhead introduced by our secure platform with public, cross-domain benchmarks.
In the remainder of the paper, Sect. 2 outlines the core concepts we used to devise our solution. Sect. 3 presents a running example that is a reference throughout the paper and outlines the problem we tackle along with the requirements it poses. In Sect. 4, we describe our approach in detail. Sect. 5 provides a formal security analysis of our solution. In Sect. 6, we present our proof-of-concept implementation and the results of the experiments we conducted therewith. In Sect. 7, we discuss the implications and limitations of our work. Sect. 8 reviews related work. Finally, Sect. 9 concludes the paper and outlines directions for future work. Following the open science principles, the source code and experiment data can be found at anonymous.4open.science/r/SPARTA.
2 Background
@@
Confidential computing Confidential computing represents a paradigm wherein sensitive data undergoes processing within a protected processing unit. Within this unit, the data and the processes treating those data remain concealed from any external entities [ConfidentialComputingBook]. Confidential computing underpins the Trusted Execution Environments (TEEs) [sabt2015trusted]. A TEE is an environment that provides a level of assurance of the following three properties: (i) data confidentiality, namely the fact that unauthorized entities cannot access or observe data while it resides within the TEE and is being processed, (ii) data integrity, namely the fact that unauthorized entities cannot modify, delete, or insert data while it remains within the TEE and is undergoing processing and (iii) code integrity, namely that unauthorized entities cannot manipulate, delete, or introduce alterations to the code executed within the TEE [confidential2022technical, Feng2024, Sardar2023]. Our work leverages Intel Software Guard Extensions (SGX) [SGX1, SGX2, SGX3, AcceleratingEncrypted], a collection of security-related instructions embedded within contemporary Intel Central Processing Unit (CPU)s. SGX facilitates the secure execution of code (deployed in the form of so-called Trusted Applications, or Trusted Apps for short) and data within enclaves, namely hardware-protected environments. The three primary security features of enclaves are isolation, attestation, and sealing. Isolation involves the enclave’s placement within a hardware-secured memory enclave page cache, specifically designated for housing protected code and data. However, TEE-allocated memory is limited. Exceeding these limits may entail stops that can severely affect availability (the service is interrupted) or data secrecy (encrypted data can be swapped to outer memory regions). Additionally, SGX enables remote attestation to allow external entities to authenticate a target enclave. During remote attestation, the external entity communicates with the attestation service to validate the integrity of the information provided by the target enclave. This process entails comparing enclave information with the trusted configuration within the target enclave, ensuring the integrity of code and data loading during the initial enclave establishment. Moreover, SGX employs sealing mechanisms to safeguard enclave contents stored beyond the enclave’s boundaries. These mechanisms entail encrypting enclave contents using a concealed sealing key, derivable from either the measurement hash or the author identity of the enclave. Both sealing mechanisms perform Advanced Encryption Standard - Galois/Counter Mode (AES-GCM) encryption. This approach guarantees exclusive access and decryption privileges to the sole corresponding enclave. Decision logic and access control One of the most well-known languages to model and formalize decision logic in business applications is Decision Model and Notation (DMN) [dmn, batoulis2015extracting, de2021integrating], a standard by the Object Management Group (OMG) 111omg.org/spec/DMN, accessed 2025-06-17 for modeling and formalizing decision logic in multi-party processes and business applications. A Decision Model and Notation (DMN), at its core, is a decision table that captures critical knowledge and automates decision logic [calvanese2016semantics]. A DMN table consists of columns representing the inputs and outputs of a decision, and rows denoting rules (the decision logic). Rules are expressed using Friendly Enough Expression Language (FEEL), a language defined within the DMN standard. Each row has an identifier, a condition for each input column, and one specific value for each output column. A rule is considered applicable when the input values satisfy all associated conditions; in that case, the corresponding output values are produced [calvanese2016semantics, de2021integrating]. To express authorization policies over the usage of resources, the Abbreviated Language for Authorization (ALFA) 222alfa.guide/alfa-authorization-language/, accessed 2025-06-17 language provides a concise, human-readable format. Built on top of the eXtensible Access Control Markup Language (XACML),333docs.oasis-open.org/xacml, accessed 2025-06-17 ALFA offers a simplified syntax for defining fine-grained, attribute-based access control (ABAC) policies based on users, actions, resources, and context. Distributed notarization and storage A blockchain is a Distributed Ledger Technology (DLT) wherein transactions are sorted, organized into blocks, and interconnected to form a chain. Blockchains possess resistance against tampering, achieved through cryptographic methods. Public blockchains like Ethereum [Ethereum] and Algorand [Algorand] permit the deployment and execution of programs named smart contracts [IntroducingEthereumandSolidity, zheng2020overview]. To mitigate the expenses associated with the invocation of smart contracts, external Peer-to-peer (P2P) systems are frequently employed for storing large volumes of data [ArchitectureforBlockchainApplications]. Among the facilitating technologies is the InterPlanetary File System (IPFS) [ipfs],444ipfs.tech, accessed 2025-07-30 a distributed system designed for file storage and access, leveraging a Distributed Hash Table (DHT) to distribute files across multiple nodes. Utilizing content-addressing, IPFS assigns a unique identifier to each file within the network. When a file is uploaded to IPFS, it generates a content-addressed resource locator based on its hash (e.g., QmbPg[…]shTg9). Changing the file changes the hash and hence makes the locator invalid. IPFS supports mutable pointers that allow content updates without changing the shared reference. Specifically, InterPlanetary Name System (IPNS) enables the creation of mutable pointers, known as IPNS names (e.g., PatientsWAggrKey: QmbPg[…]shTg9). These names point to the content-addressed locators, and can be updated with newer locators while retaining the verifiability of content addressing.
3 Problem illustration, running example, and requirements
@@
Table I: An excerpt of the data fields provided to the system by the process participants of Sect. 3
@@
To illustrate the problem we address, we introduce a real-world scenario from the healthcare domain: a vaccine distribution campaign inspired by [vaccinedistribution, noh2021group] and linked to a research project funded by the European NextGenerationEU. [^5] To comply with the double-blindness requirement of the conference, we will reveal the project only upon acceptance. We use this scenario as a running example and motivating case study throughout the paper.
In our scenario, a central medical hub (henceforth, medical hub for short) receives vaccine supplies and handles their dispatch in a federal State. The downstream process involves other participants, namely patients who get an injection, vaccination centers who administer the injections, and several specialized carriers delivering the vaccines on premise. It covers two stages. In the first phase, the involved participants act as Data Providers by contributing data into the system. We will refer to it as the data provision stage. In our scenario, e.g., the medical hub reports received vaccine stocks; patients submit vaccine requests including personal information; vaccination centers update storage capacity and vaccination progress; and carriers register their service conditions. Table I contains part of the data these participants must provide to the system. In a subsequent stage (data usage), authorized participants retrieve and read the data provided earlier, either specific to single entities or aggregated, and receive decision support depending on their role in the process. For example, patients can retrieve their declarations at any time, carriers can be informed about the on-time delivery rate, and the medical hub requests automated support for vaccine supply distribution. Decisions related to the vaccine distribution include (i) the priority to be assigned to patients to administer the vaccine dose (which we will use as a running example throughout this paper), (ii) the restocking need for each vaccination center, (iii) and which carrier should handle each delivery.
In our example, several critical aspects need to be considered. First, all data is confidential and must be protected from unauthorized access. For example, revealing a single patient’s metadata constitutes a privacy breach. Second, data integrity and availability are critical. Any unauthorized modification could alter decisions or falsify patient records, while system failures or data unavailability could disrupt the decision-making process. Ensuring privacy, integrity, and availability while enabling a customizable, secure and verifiable automated decision-making is an open challenge motivating our research.
@@
4 Approach
In this section, we present our approach named Secure Platform for Automated decision Rules via Trusted Applications (SPARTA). We first introduce the system components (Sect. 4.1), and then describe how they interact to manage information and enable secure decision-making over confidential data (Sect. 4.2).
4.1 Actors and basic components
@@
Figure 1 provides a bird’s eye view of the SPARTA approach. We identify four main actors, each responsible for a distinct functionality. The Policymaker defines the decision rules and specifies which functions are available, along with the participant attributes or attribute combinations required to invoke them. Considering the example in Sect. 3, the Policymaker role can be fulfilled either by a regulatory body such as the Ministry of Health, or by the consortium directly involved in the decision process (upon consensus, as in [iacr/CerulliCNPS23]). The Data Provider provides the input data and contextual information used for decision-making. In our example, the Data Providers are the participants: the medical hub, patients, vaccination centers, and the carriers. The Certification Authority (CA) signs a Certificate Signing Request (CSR) generated by a Registration Authority (RA), thereby producing a valid certificate. The RA handles participant registration and verifies their identity and role. Once verification is complete, it generates the CSR. By signing the CSR, the CA attests actor metadata, authenticates participants, and certifies their roles and attributes, determining their eligibility to interact with the system and invoke specific functions. In our example, the role of the CA should be fulfilled by a globally recognized Certification Authority (e.g., DigiCert, Entrust, or GlobalSign), [^6] DigiCert: digicert.com; Entrust: entrust.com; GlobalSign: globalsign.com/en, accessed 2025-08-04 akin to what is offered by the National Health Service in the United Kingdom [DBLP:journals/istr/Dalton03]. The RA role should be covered by a National Health Service Provider or a government eID system for patients; a Regional Department of Public Health or Healthcare Ministry for vaccination centers; and a Ministry of Transport or logistics regulator for carriers. Finally, the Decider receives decision support based on the available data. If permitted by access policies, the Decider may also retrieve raw data and invoke aggregation functions. A Decider is, e.g., the medical hub.
@@
The interplay of five basic components realizes the architecture. (I) The Confidential Computing Unit (CCU) is SPARTA’s core component. We deploy the CCU inside a Trusted Execution Environment (TEE). This hardware-secured environment guarantees data confidentiality and ensures the integrity of both computation and code (see Sect. 2), thereby meeting Tab. II. Notice that we make CCU replicable for the sake of information availability and system robustness (as per Tabs. II and II; we explain how we do so later in Sect. 4.2.1). The CCU deploys a separate Decision Trusted App (DeTApp) for each decision process in a dedicated enclave within the TEE. A DeTApp exposes an invocable Decision Support Object (DeSObj) providing and controlling access to the Policymaker’s decision-making functions. The DeSObj, in turn, calls a Decision Function Object (DeFObj) realizing those functions, and a linked Decision Library (DeLib), a predefined library that offers utilities for DeSObjs such as data retrieval and decryption, policy checking, and certificate validation. DeSObjs and DeFObjs are customizable and are built during SPARTA’s initialization phase (see Sect. 4.2.1 next) by a DeSObj builder and a DeFObj builder, respectively, based on a policy descriptor and a decision model provided by the Policymaker. A DeTApp builder binds those objects together with the DeLib and deploys the DeTApp in the enclave.
@@
(III) The Admin Interface handles information exchange with CA and Policymaker. (IV) The User Interface is the means through which the Data Providers and the Decider interact with the CCU. Finally, SPARTA is paired with (V) the Data Manager, an external component providing two architectural buttresses: (i) The Data File Processor, storing the content of the exchanged data, and (ii) The Notarization System, saving and sealing the resource locators to those data. To meet Tab. II, we leverage two technologies for the realization of the above components. We employ IPNS over IPFS for the Data File Processor because IPNS guarantees permanence, immutability and content-addressing for stored files (see Sect. 2). We use a public blockchain platform as the Notarization System to store the IPNS names in a tamper-proof, available and traceable way. Other solutions like STORJ777storj.io, accessed 2025-07-30 and distributed relational databases [distributed-relational-database] can be used, respectively, in scenarios with less restrictive security requirements.
Next, we elucidate the interplay of these components.
4.2 Workflow
The workflow of SPARTA mainly consists of two phases, as illustrated in Fig. 1: First, the system initialization (Sect. 4.2.1) prepares the necessary data structures and programs, then the execution follows (Sect. 4.2.2). SPARTA’s workflow involves several cryptographic keys. Table VI summarizes these keys along with the Sections in which they play a crucial role, helping the reader better understand their usage throughout the workflow.
4.2.1 System initialization